# MentalMAC: Enhancing Large Language Models for Detecting Mental Manipulation via Multi-Task Anti-Curriculum Distillation

### v1

@@ A.1 Example Demonstration
Our proposed EvoSA is a prompting method for synthesizing new dialogues by combining two parent dialogues into a single child dialogue. An example of such synthesis from MentalManip is shown in Figure 4. The blocks of the same colour represent portions of the child dialogue that closely resemble the parent dialogue. As observed, similar to the evolutionary algorithm, the child dialogue incorporates elements from the parent dialogue. Furthermore, the child dialogues exhibit a rich variety of speech acts, distinguishing them from the parent dialogues in terms of background, roles, and content. In addition, as illustrated in Figure 5, we found that EvoSA can extend a two-person dialogue to a multi-person conversation, demonstrating its potential to synthesize dialogues involving multiple participants.
In data synthesis, we did not find refusals to answer on MentalManip, probably because MentalManip does not contain excessively toxic text. However, during the expansion of ReaMent, we observed that the LLM declined to respond in 257 out of 3,257 instances as a result of its safety mechanisms. Although there is less than a 10% chance of refusal to respond, this does not significantly affect our ability to generate sufficient data, provided we have a certain number of parent dialogues.
@@ A.2 Quality of Child Dialogues
In selecting evaluators, we recruited candidates via a voluntary application process. They were informed about the potential exposure to offensive content and provided their consent. Subsequently, the candidates received training on mental manipulation. After that, we randomly selected 100 dialogues from MentalManipcon for candidates to annotate. Only those with 80% annotation accuracy were considered as evaluators. In the end, we recruited 3 evaluators to evaluate the quality and label consistency of the child dialogues.
@@
Please rate the quality of this dialogue on a 5-point scale (1 = very poor, 5 = excellent), taking into account factors such as coherence, logical consistency, and naturalness.
@@
We collected the scores assigned by the three evaluators to the Yes and No labels, and reported their average values in Table 6. Analysis shows that child dialogues labeled No tend to be of lower quality than those labeled Yes. On average, child dialogues labeled Yes and No received scores of 4.265 and 3.788, respectively. The overall mean score of 4.022 indicates that EvoSA effectively guides LLMs to generate higher-quality dialogues. Figure 6 illustrates the correlation between dialogue length and quality score. The regression line shows that as dialogue length increases, quality tends to improve, likely due to the additional context provided by longer dialogues. The Pearson correlation coefficient (r = 0.270, p = 0.0004) indicates a statistically significant positive correlation, suggesting that dialogue length does have an effect on the quality score in EvoSA-generated dialogues.
@@ A.3 Label Consistency Between Child and Parent Dialogues
To assess label consistency, evaluators were asked to analyze why parent dialogues were labeled as containing or not containing mental manipulation. Based on their analysis, they then determined whether the labels of the child dialogues were consistent with those of the parent dialogues. We provided the following instruction for evaluators to complete the task:
@@
Table 7 presents the results of the labeling consistency assessment. When two out of three evaluators consider the labels consistent ("$\geq 2$"), we regard the child dialogue as consistent with the parent dialogues in terms of label. Of the child dialogues labeled No and Yes, 89.2% and 96.9%, respectively, are being consistent with the parent dialogue labels. This indicates that dialogues without mental manipulation are being rated lower than those containing mental manipulation, likely because LLMs are tending to mistakenly classify non-manipulative dialogues as manipulative.
@@
Table 7: Results of the label consistency assessment. "$<2$" and "$\geq 2$" denote that fewer than 2 and at least 2 people, respectively, believe the labels are consistent.
@@
where $c_{i}$ denotes the label consistency score of the $i$-th child dialogue; $e_{ji}$ denotes the $j$-th evaluator’s rating of the $i$-th child dialogue in terms of label consistency, taking the value 1 if consistent and 0 otherwise; and $n_{e}$ denotes the number of evaluators. Figures 7 and 8 illustrate the relationships between label consistency scores and dialogue length, and dialogue quality, respectively. The Pearson correlation coefficient indicates that there is no significant relationship between label consistency scores and dialogue length, but it shows a significant positive correlation with dialogue quality.
@@ B.1 Prompt Templates for Teacher Models
where <does not contain/contains> is given based on true labels. The definition of mental manipulation is adopted from Wang et al. (2024). As can be seen, we require the teacher model not only to reason step by step, but also to provide detailed evidence explaining why the dialogue does or does not conform to known manipulation tactics. In order to generate incorrect answers, here we prompt the teacher model with incorrect labels.
@@ B.2 Prompt Templates for Student Models
Obviously, the student model prompt could not mirror the teacher model prompt exactly; therefore, we designed a simplified version specifically for the student. The prompt template for Task 1 is
@@
The prompt template for Task 2 makes the student model first output a binary judgment (Yes or No). To save computational resources, we use the same prompt for both Task 1 and Task 2. For Task 1, we compute the loss solely on the token corresponding to the binary judgement.
@@ C.1 Details on the Annotation Process
In selecting annotators, we recruited candidates via a voluntary application process. They were informed about the potential exposure to offensive content and provided their consent. Subsequently, the candidates received training on mental manipulation. After that, we randomly selected 100 dialogues from MentalManipcon for candidates to annotate. Only those with 80% annotation accuracy were considered as annotators. In the end, we recruited 12 annotators and divided them into groups of three each annotating the same data.
@@
Based on the definition of mental manipulation, please determine whether the given dialogue contains elements of mental manipulation. If it does, label it as 1; otherwise, label it as 0. In addition, provide your confidence level in the annotation on a 5-point scale (1 = very uncertain, 5 = very confident).
@@
where the definition of mental manipulation comes from Wang et al. (2024). Following this instruction, we collected both the labels assigned by all annotators and their corresponding confidence levels.
@@ C.2 Detailed Descriptions
Table 8 presents the distribution of binary labels. As shown, the number of dialogues labeled as No is slightly smaller than those labeled as Yes.
@@ C.3 Analysis of Annotation Quality
Figure 10: Distribution of label confidence scores. "Mean" refers to the average value, while "Q1" and "Q2" is the first quartile and the median, respectively.
@@ D.1 Parameters for Teacher Model
The teacher model performs two tasks: the first involves the unsupervised synthesis of new dialogues using EvoSA, and the second concerns the construction of multitask data. For both tasks, we use identical parameters, setting the temperature to 0 and the maximum token limit to 512.
@@ D.2 Parameters for Student Model
During training, the student model is trained with a maximum sequence length of 1500 tokens. In the first two stages we train 2 epochs, and in the last stage we train 2 epochs on MentalManip and 1 epoch on ReaMent. In each stage, learning rate warm-up with a constant schedule is applied, with the warm-up ratio set to 0.05.

### v2

@@ A.1 Example Demonstration
Our proposed EvoSA is a prompting method for synthesizing new dialogues by combining two parent dialogues into a single child dialogue. An example of such synthesis from MentalManip is shown in Figure 4. The blocks of the same colour represent portions of the child dialogue that closely resemble the parent dialogue. As observed, similar to the evolutionary algorithm, the child dialogue incorporates elements from the parent dialogue. Furthermore, the child dialogues exhibit various speech acts, distinguishing them from the parent dialogues regarding background, roles, and content. In addition, as illustrated in Figure 5, we found that EvoSA can extend a two-person dialogue to a multi-person conversation, demonstrating its potential to synthesize dialogues involving multiple participants.
In data synthesis, we did not find refusals to answer on MentalManip, probably because MentalManip does not contain excessively toxic text. However, during the expansion of ReaMent, we observed that the LLM declined to respond in 257 out of 3,257 instances due to its safety mechanisms. Although there is less than a 10% chance of refusal to respond, this does not significantly affect our ability to generate sufficient data, provided we have a certain number of parent dialogues.
@@ A.2 Quality of Child Dialogues
In selecting evaluators, we recruited candidates via a voluntary application process. They were informed about the potential exposure to offensive content and consented. Subsequently, the candidates received training on mental manipulation. Afterward, we randomly selected 100 dialogues from MentalManipcon for candidates to annotate. Only those with 80% annotation accuracy were considered as evaluators. In the end, we recruited three evaluators to evaluate the quality and label consistency of the child dialogues.
@@
Please rate the quality of this dialogue on a 5-point scale (1 = very poor, 5 = excellent), considering factors such as coherence, logical consistency, and naturalness.
@@
We collected the scores assigned by the three evaluators to the Yes and No labels, and reported their average values in Table 6. Analysis shows that child dialogues labeled No are of lower quality than those labeled Yes. On average, child dialogues labeled Yes and No received scores of 4.265 and 3.788, respectively. The overall mean score of 4.022 indicates that EvoSA effectively guides LLMs to generate higher-quality dialogues. Figure 6 illustrates the correlation between dialogue length and quality score. The regression line shows that as dialogue length increases, quality tends to improve, likely due to the additional context provided by more extended dialogues. The Pearson correlation coefficient (r = 0.270, p = 0.0004) indicates a statistically significant positive correlation, suggesting that dialogue length affects the quality score in EvoSA-generated dialogues.
@@ A.3 Label Consistency Between Child and Parent Dialogues
To assess label consistency, evaluators were asked to analyze why parent dialogues were labeled as containing or not containing mental manipulation. Based on their analysis, they then determined whether the labels of the child dialogues were consistent with those of the parent dialogues. We provided the following instructions for evaluators to complete the task:
@@
Table 7 presents the results of the labeling consistency assessment. When two out of three evaluators consider the labels consistent ("$\geq 2$"), we regard the child dialogue as consistent with the parent dialogues in terms of label. Of the child dialogues labeled No and Yes, 89.2% and 96.9%, respectively, are consistent with the parent dialogue labels. This indicates that dialogues without mental manipulation are being rated lower than those containing mental manipulation, likely because LLMs are tending to classify non-manipulative dialogues as manipulative mistakenly.
@@
Table 7: Results of the label consistency assessment. "$<2$" and "$\geq 2$" denote that fewer than two and at least two people, respectively, believe the labels are consistent.
@@
where $c_{i}$ denotes the label consistency score of the $i$-th child dialogue; $e_{ji}$ denotes the $j$-th evaluator’s rating of the $i$-th child dialogue in terms of label consistency, taking the value 1 if consistent and 0 otherwise; and $n_{e}$ denotes the number of evaluators. Figures 7 and 8 illustrate the relationships between label consistency scores and dialogue length, and dialogue quality, respectively. The Pearson correlation coefficient indicates no significant relationship between label consistency scores and dialogue length, but shows a significant positive correlation with dialogue quality.
@@ B.1 Prompt Templates for Teacher Models
where <does not contain/contains> is given based on true labels. The definition of mental manipulation is adopted from Wang et al. (2024). As can be seen, we require the teacher model not only to reason step by step, but also to provide detailed evidence explaining why the dialogue does or does not conform to known manipulation tactics. Here, we prompt the teacher model with incorrect labels to generate incorrect answers.
@@ B.2 Prompt Templates for Student Models
The student model prompt could not precisely mirror the teacher model prompt; therefore, we designed a simplified version for the student. The prompt template for Task 1 is
@@
The prompt template for Task 2 makes the student model first output a binary judgment (Yes or No). We use the same prompt for both Task 1 and Task 2 to save computational resources. For Task 1, we compute the loss solely on the token corresponding to the binary judgement.
@@ C.1 Details on the Annotation Process
In selecting annotators, we recruited candidates via a voluntary application process. They were informed about the potential exposure to offensive content and consented. Subsequently, the candidates received training on mental manipulation. Afterward, we randomly selected 100 dialogues from MentalManipcon for candidates to annotate. Only those with 80% annotation accuracy were considered as annotators. Ultimately, we recruited 12 annotators and divided them into three groups, annotating the same data.
@@
Based on the definition of mental manipulation, please determine whether the given dialogue contains elements of mental manipulation. If it does, label it 1; otherwise, label it 0. In addition, provide your confidence level in the annotation on a 5-point scale (1 = very uncertain, 5 = very confident).
@@
where the definition of mental manipulation comes from Wang et al. (2024). Following this instruction, we collected the labels that all annotators assigned and their corresponding confidence levels.
@@ C.2 Detailed Descriptions
Table 8 presents the distribution of binary labels. The number of dialogues labeled as No is slightly smaller than those labeled as Yes.
@@ C.3 Analysis of Annotation Quality
Figure 10: Distribution of label confidence scores. "Mean" refers to the average value, while "Q1" and "Q2" are the first quartile and the median, respectively.
@@ D.1 Parameters for Teacher Model
The teacher model performs two tasks: the unsupervised synthesis of new dialogues using EvoSA, and the construction of multitask data. We use identical parameters for both tasks, setting the temperature to 0 and the maximum token limit to 512.
@@ D.2 Parameters for Student Model
During training, the student model is trained with a maximum sequence length of 1500 tokens. In the first two stages, we train two epochs; in the last stage, we train two epochs on MentalManip and one on ReaMent. A learning rate warm-up with a constant schedule is applied in each stage, with the warm-up ratio set to 0.05.